OpenStudy (anonymous):

What value is added to both sides of the equation x2 − 4x = −2 in order to solve by completing the square

OpenStudy (mathstudent55):

Take the coefficient of the x term. Divide it by 2. Then square it.

OpenStudy (anonymous):

okay so it would be 4

OpenStudy (malikasif):

yeah you divide -4 by 2

OpenStudy (malikasif):

it would be -2 then adding -2 square on both sides for completing square

OpenStudy (wolf1728):

So we take the equation x2 − 4x = −2 then add 4 to both sides x2 − 4x +4 = −2 +4 Now we take the square root of both sides x-2 = sq root (2)